OpenStudy (anonymous):

Question about 2nd-derivative test for finding relative max/min (posting below)

OpenStudy (anonymous):

I need to find the relative max/min values for this function: \[g(x) = x + 2/\sqrt{x}\] I found the 1st derivative which is g'(x) = 1 + (-x^-3/2) I found the critical numbers which are 0 and 1 Then the 2nd derivative, which is g''(x) = (3/2)x^(-5/2) g''(0) = 0 g''(1) > 0 (1,f(1)) is a relative minimum But what does it mean when g''(0) = 0?

OpenStudy (anonymous):

Wait ya stupid mistake. 0 is not a critical number

OpenStudy (anonymous):

sorry for my silly mistake:p

OpenStudy (aum):

Well, you realized 0 is not a critical number. Also, g''(0) is not zero but undefined.

OpenStudy (imstuck):

Not to butt in, but how is 0 not a critical number? When you set the second derivative = to 0, then x = 0, doesn't it? @aum

OpenStudy (aum):

x = 0 is not in the domain of g(x). g(x) is only defined for x > 0 \(\Large g''(x) = \frac 32 \frac{1}{x^{5/2}}\) So g''(0) is undefined.

OpenStudy (aum):

We say that x = c is a critical point of the function f(x), IF f(c) exists AND either f'(c) = 0 OR f'(c) does not exist. Note that we require that f(c) exists in order for x = c to actually be a critical point. This is an important, and often overlooked, point. Here g(0) does not exist and therefore x = 0 is not a critical point.